OpenStudy (anonymous):

Find the equation of the perpendicular line to each of the following that passes through the point indicated 4x + 5y = - 40 passing through the point (-1, 3)

OpenStudy (anonymous):

that isn't one of my answer choices

OpenStudy (anonymous):

To do this sort of question, I'd first get the answer in terms of y=mx+c. Then m is your gradient and to get the perpendicular gradient it is -1/m. Then you need to get the y-intercept right so that it goes through that point! To do this, replace x with -1 in your y=mx equation then you should be able to work out what c has to be so that in that instance, y is 3! Does that make sense?

OpenStudy (anonymous):

y=5/4x+17/5

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

*17/4

OpenStudy (anonymous):

5x-4y = -17

OpenStudy (anonymous):

There you go.

OpenStudy (anonymous):

-4/5 x - 8

OpenStudy (anonymous):

Ok, so you can rewrite that equation as y=−4/5x−8 agreed?